solve the simultaneous equations:
x2 + y2 = 16
y = 3x -1
please give answers to an accuracy of 2 decimal places

Answer by scott8148(6628)

substituting ___ x^2 + (3x - 1)^2 = 16

x^2 + 9x^2 - 6x + 1 = 16 ___ 10x^2 - 6x - 15 = 0

use the quadratic formula to find the two values of x , then substitute to find the two values of y